OpenStudy (anonymous):

~~~~~~~~~~~~~FAN AND MEDAL~~~~~~~~~~~~~~ For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (anonymous):

you go to a (cheap) bar 2 shots and a beer cost $10 3 shots and 4 beers cost $25 how much is a beer, and how much is a shot?

OpenStudy (anonymous):

if your teacher will not appreciate beers and shots, turn them in to hot dogs and sodas

OpenStudy (anonymous):

lmaoo Okay but i dont know how to put this all together

OpenStudy (anonymous):

you mean how do you solve it?

OpenStudy (anonymous):

Oh nevermind This is pretty easy can you just come back when I'm finished so you can tell me if It looks right? :D

OpenStudy (anonymous):

\[2x + y = 10 \\ 3x + 4y = 25\] first equation is the same as \(y=-2x+10\) so you can solve \[3x+4(-2x+10)=25\] for \(x\) then solve for \(y\)

OpenStudy (anonymous):

Can you teach me how to solve for x , and y please

OpenStudy (anonymous):

\[3x+4(-2x+10)=25\] multiply out using the distributive law \[3x-8x+40=25\] collect like terms\[-5x+40=25\] subtract \(40\) \[-5x=-15\] divide by \(-5\) \[x=3\]

OpenStudy (anonymous):

Ok thank you and what about the y

OpenStudy (anonymous):

\[2x+y=10\\ 2\times 3+y=10\]is a start

OpenStudy (anonymous):

Lol then its 4 eh